OpenStudy (anonymous):

If a bateria culture starts with 4000 bateria and doubles every 30 minutes, how many minutes will it take the population to reach 42000?

OpenStudy (anonymous):

\[4000\times 2^{\frac{t}{30}}=42000\] solve for t

OpenStudy (anonymous):

can you finish from here?

OpenStudy (anonymous):

um sorta no.... I really suck at exponets... Im so dead for the final Friday How do I do it

OpenStudy (anonymous):

Don't tell me! Just tell me how to solve. No answering

OpenStudy (anonymous):

start by dividing by 4000

OpenStudy (anonymous):

then you can solve for \[\frac{t}{30}\] in one step if you recall to solve \[b^x=A\] for x you write \[x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

10.5= 2^(t/30)

OpenStudy (anonymous):

ok then one more step to get \[\frac{t}{30}\]

OpenStudy (anonymous):

divide by 2 on both sides?

OpenStudy (anonymous):

no, now you have to use logarithms

OpenStudy (anonymous):

oh... this is where i mess up T.T um ok 1n(10.5)/1n (2) =t/30

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so \[t=\frac{30\ln(10.5)}{\ln(2)}\] and now a calculator. will get the answer in minutes i mean the answer will have minutes for the units

OpenStudy (anonymous):

i got 3.3923174 = t/30 so multiply 30 on both sides and I got 101.769

OpenStudy (anonymous):

Omg thanks so mcuh! I def. Love this site I hope I can help out too just like you helped me <3

OpenStudy (anonymous):

that is what i get too http://www.wolframalpha.com/input/?i=30ln%2810.5%29%2Fln%282%29

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

XD that site ive been trying to avoid using...

OpenStudy (anonymous):

well i just use it as a calculator. we can do your last problem too if you like. i think it is easier than you think

OpenStudy (anonymous):

Oh? Ok! Which was that I forgot already